OpenStudy (kg1975):

Plz Help @Taylor0402

OpenStudy (taylor0402):

go on

OpenStudy (kg1975):

that made no sence at all! to me

OpenStudy (kg1975):

how did u get 1 as the answer?

OpenStudy (taylor0402):

ok can you put the question back in so i can help you understand without having to switch back and forth between questions please

OpenStudy (kg1975):

What is the value of the expression \[\frac{ 4x-y }{ 2y+x }\] when x = 3 and y = 3?

OpenStudy (taylor0402):

ok this is quite simple so we know we have to substuite x for 3 and y for 3 also so plug in x and y on the equation\[4 \times3-3 \over 2\times3+3\]

OpenStudy (taylor0402):

you got that so far

OpenStudy (kg1975):

yes

OpenStudy (taylor0402):

do you understand why we are multiplying 4x and 2y to get 4 times 3 and 2 times 3

OpenStudy (kg1975):

uhm no

OpenStudy (taylor0402):

ohhh ok now i see the problem

OpenStudy (kg1975):

sorry :( i always fail math

OpenStudy (kg1975):

im horrible at it DX

OpenStudy (taylor0402):

no no its ok i just learned this in 7th grade so its easier for me to understand im a freshman now

OpenStudy (kg1975):

im in 10th grade now and i still dont understand it DX

OpenStudy (unofficialllyy):

Try doing it in the order of PEMDAS. P=Parenthesis E=Exponents M=Multiplication D=Division A=Addition S=Subtraction So you plug in 3 for x, and 3 for y. \[\frac{ 4(3)-3 }{ 2(3)+3}\]

OpenStudy (taylor0402):

but anyways if any number (for example we will use 3) has a variaple next to it we always multiply unless it says to add subtract or divide. So 3m were m=4 is 3 times 4. but if were were to do 3+m it would be 3+7=7

OpenStudy (unofficialllyy):

So on the top you would do 4 times 3 to get what?

OpenStudy (taylor0402):

@Unofficialllyy you can go i got this no need to confuse them

OpenStudy (kg1975):

???D: im very confused

OpenStudy (unofficialllyy):

Just look at this way. On the top you have 4 multiplied by 3. Do you know what that is?

OpenStudy (taylor0402):

thats ok so if any number has a variable next to it (a letter) then you ALWAYS multiply unless it says to do otherwise

OpenStudy (taylor0402):

@Unofficialllyy he dosent know why he is multiplying those two numbers please go away

OpenStudy (kg1975):

she*

OpenStudy (unofficialllyy):

You're not explaining it correctly though.

OpenStudy (taylor0402):

sorry @kg1975

OpenStudy (kg1975):

im gonna make a new chat. for you ok taylor

OpenStudy (taylor0402):

ok thanks unofficialllyy has no idea

OpenStudy (taylor0402):

dude will you just leave

OpenStudy (unofficialllyy):

No need to be rude, just trying to help. That's what this site is for! :)

OpenStudy (kg1975):

i tagged u in a room

OpenStudy (taylor0402):

yea but i have this so go find another question ;)

OpenStudy (unofficialllyy):

I won't try to help anymore, I'm just going to spectate. :)